I can't figure out this problem which is part of the Pythagarean Therum worksheet:
From Canoeville it is 2.4 kilometers to White Beach and 3.0 kilometers to the Lodge. How far is it from White Beach across the lake to the Lodge?

Answer by checkley75(3666)

TO APPLY THE PAYTHAGAREAN THERUM WE NEED TO WHICH IS THE HYPOTENUSE(LONGEST SIDE). YOU HAVEN'T INDICATED WHICH DISTANCE IS THE LONGEST.
2.4^2+x^2=3^2
5.76+x^2=9
x^2=9-5.76
x^2=3.24
x=1.8 km.